Fermi National Laboratory
 

Accelerator Update for May 16, 2007 - May 18, 2007

 

Wednesday May 16

The day shift began with the Tevatron (TeV) in access, with Operations monitoring the Recycler stash (250.1E10), with Antiproton Source (Pbar) experts conducting reverse proton studies, with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) in access.

At 8:34 AM, a Linac RF station (LRF3) tripped off and wouldn’t reset. An expert investigated and had it back up and running ten minutes later.

At 9:04 AM, LRF3 tripped off again. Operators and experts investigated. Later, experts reported that the voltage setting for its power amplifier had drifted high.

At 10:08 AM, LRF3 tripped off. Operators investigated and had it back online within 15 minutes.

LRF3 was back on by 10:48 AM.

MI experts began some studies at 11:08 AM.

At 2:22 PM, LRF3 tripped off. It reset.

Pbar resumed stacking at 3:24 PM.

At 3:50 PM, LRF3 tripped off. It reset.

Operations reported at 3:59 PM that all the TeV keys had been returned, Pbar was conducting stacking studies, NuMI was taking beam, and that CDF and D0 were still in on access.

D0 came out of access at 7 PM.

An MI expert completed his MI-8 collimator study at 7:16 PM.

Pbar experts conducted their first transfer of antiprotons to the Recycler with their new helix at 8:26 PM.

At 10:46 PM, LRF3 tripped off. It reset.

Operations reported at 11:54 PM that the Recycler had stashed four times during the evening shift.

Thursday May 17

The midnight shift began with Operations monitoring the stash (267.8E10), with Pbar stacking (9.7E10), and with NuMI and MiniBooNE taking beam.

Operations reported at 12:18 AM, that TeV experts had replaced a loop controller. The failed controller was inhibiting the dump loop.

Operations put the TeV through a dry squeeze at 12:38 AM.

At 1:33 AM, a MI RF station (MIRF14) tripped off and wouldn’t reset. Operators bypassed the station.

At 5:23 AM, Operations prepared the TeV for an F sector and Transfer hall access.

At 5:50 AM, LRF3 tripped off. Operators reset the station.

The TeV F sector access ended at 7:41 AM.

Operations reestablished beam to MiniBooNE at 7:46 AM.

Operations reported at 8:09 AM that the Recycler had stashed twice during the midnight shift.

MI RF experts brought MIRF14 back online at 8:26 AM.

Operations resumed sending beam to Pbar and NuMI at 8:55 AM.

At 8:58 AM, stacking stopped due to a Pbar Debuncher power supply (D:H704) tripping off. Operators had the supply reset by 9:18 AM. Stacking resumed.

At 9:43 AM, LRF3 tripped off. Experts investigated. They had the station back on by 10:20 AM.

At 10:26 AM, LRF3 tripped off. Experts investigated and determined that its Power Amplifier (PA) needed to be replaced.

At 10:43 AM, Operations reported that the Recycler lost half of its stash while experts were investigating a problem with a power supply (QCL). Also, TeV experts found a ground fault on a TeV power supply (T:SD). Later, TeV experts traced the problem to the ground fault detector circuit.

Operations reported at 10:43 AM that they had switched to an energy conservation time line in the Time Line Generator (TLG) for the duration of the LRF3 work.

Operations reported at 10:53 AM that the Run Coordinator would allow some maintenance work during the LRF3 repairs.

At 11:54 AM, Operations reported that the Booster west anode power supply (WAPS) tripped off. It was back online in 10 minutes.

Operations put the TeV through a dry squeeze at 12:32 PM.

The LRF3 work was completed by 1:35 PM. Linac resumed sending beam at 1:36 PM.

Pbar resumed stacking at 2:34 PM.

Operations put the TeV through a wet squeeze at 2:56 PM.

At 3:20 PM, Operations put NuMI into a “beam to NuMI only mode” while Pbar experts checked out a problem with an Accumulator RF station (ARF1).

Pbar resumed stacking at 3:33 PM.

Operations reported at 3:56 PM that the Recycler had stashed once during the day shift.

Operations put the TeV into shot setup around 4 PM.

At 4:14 PM, LRF3 tripped off. It reset.

At 4:25 PM, the P1 transfer line beam permit tripped off due to the current on a quadrupole power supply drifting high. This also caused NuMI to trip off. Everything reset okay.

At 4:56 PM, the Booster beam valve (VI-Q5) indicated that it was moving. This caused the Booster chopper to (B:CHOP) to trip off, which halted all beam as well as shot setup. Operators and Booster experts investigated; they verified that the valve had closed. At 5:44 PM, Operations prepared Linac, the Booster chopper is located in the downstream area of Linac, for an access.

Booster experts reported 6:08 PM that they found a broken airline for VI-Q5.

Operations reported that operators and Booster experts had repaired the airline for VI-Q5. Beam resumed to NuMI, Pbar, and MiniBooNE around 6:40 PM. Shot setup also resumed.

Operations established store 5435 at 8:15 PM with an initial luminosity of 142.99E30.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Friday May 16

The midnight shift began with Operations monitoring store 5435 (84.08E30) and the stash (88.0E10), with Pbar stacking (15.7E10), and with NuMI and MiniBooNE taking beam.

At 5:36 AM, LRF3 tripped off.

Operations reported at 7:58 AM that the Recycler had stashed four times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator may allow Pbar, Recycler, and MI experts to conduct studies that have little impact on stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/18/2007    email Fermilab